OpenStudy (anonymous):

Part 1: Explain, in complete sentences, which method you would use to solve the following system of equations. (1 point) Part 2: Explain why you chose that method (1 point) Part 3: Provide the solution to the system. (2 points) x – 2y + z = 0 2x – 3y – 4z = –9 x + 2y – 5z = 0 1. I used elimination method 2. I used it cause it's easier than the rest. 3. I got ((97/8,10/3,23/18) Can anyone double-check my solution?

OpenStudy (anonymous):

@waterineyes

OpenStudy (anonymous):

Sure why not..

OpenStudy (anonymous):

Yeah I will also prefer elimination here..

OpenStudy (anonymous):

Reason is: The first and third equation consists same x and 2y so that they can be cancelled out by just addition or subtraction..

OpenStudy (anonymous):

Let me solve this please wait.

OpenStudy (anonymous):

K no problem :)

OpenStudy (anonymous):

Add first and third equation you will get: 2x - 4z = 0 2x = 4z ---------------------1 Now multiply first equation with 3 and second equation by 2: 3x – 6y + 3z = 0 4x – 6y – 8z = –18 Now subtract them: -x + 11z = 18 -2z + 11z = 18 9z = 18 z = 2

OpenStudy (anonymous):

I got z = 2 check your solution once again @Loujoelou

OpenStudy (anonymous):

K. I thought my solution was a bit weird.

OpenStudy (anonymous):

K I see how you got z, and I realize x=4 correct?

OpenStudy (anonymous):

@waterineyes Now I got (4,3,2)

OpenStudy (anonymous):

x = 2z if z = 2 then surely x will be 4..

OpenStudy (anonymous):

Yes you are right now..

OpenStudy (anonymous):

May I suggest you something?

OpenStudy (anonymous):

Of course.

OpenStudy (anonymous):

You can verify your answer.. See you got 4, 3 and 2, Just plug in these values in the equation given and check whether they are coming equal.. Like for first: x – 2y + z = 0 4 - 2(3) + 2 = 0 6 - 6 = 0 0 = 0 This means your answer is correct.. Now similarly put in second and third you will get 0 in each case.. This is how we verify our answers..

OpenStudy (anonymous):

K :D Tyvm @waterineyes

OpenStudy (anonymous):

Welcome Dear.. \(\huge \checkmark\)